### v4

# Restriction of $p$-modular representations of $U(2,1)$ to a Borel subgroup

Abstract
[…]
1 Introduction
Let $G$ be the unitary group $U(2,1)(E/F)$ defined over a non-archimedean local field $F$ of odd residue characteristic $p$, and let $B$ be the standard Borel subgroup of $G$. In this paper, we investigate the restriction of irreducible smooth $\overline{\mathbf{F}}_{p}$-representations of $G$ to $B$.
Our first main result is concerning principal series of $G$:
Theorem 1.1.
(Corollary 4.5, 4.6) Let $\pi$ be a smooth representation of $G$. We have:
$(1)$.  Let $\varepsilon$ be a character of $B$ such that $\varepsilon\neq\eta\circ\textnormal{det}$ for any character $\eta$ of $E^{\times}$. Then,
$\textnormal{Hom}_{B}(\textnormal{ind}^{G}_{B}\varepsilon,\pi)\cong\textnormal{Hom}_{G}(\textnormal{ind}^{G}_{B}\varepsilon,\pi)$.
$(2)$. For the trivial character of $B$, we have
$\textnormal{Hom}_{G}(\textnormal{ind}^{G}_{B}1,\pi)\cong\textnormal{Hom}_{B}(St,\pi)$.
Here $St$ is the Steinberg representation of $G$.
[…]
The following is our main result on supersingular representations of $G$:
Theorem 1.2.
We have:
$(1)$. (Theorem 5.8)  Let $\pi$ be a supersingular representation of $G$. Then
$\pi\mid_{B}$ is irreducible.
$(2)$. (Theorem 5.10) Give two smooth representations $\pi$ and $\pi^{\prime}$ of $G$. Suppose $\pi$ is supersingular. Then, we have
$\textnormal{Hom}_{G}(\pi,\pi^{\prime})\cong\textnormal{Hom}_{B}(\pi,\pi^{\prime}).$
An immediate application of $(1)$ above gives that the usual Jacquet module of a supersingular representation vanishes (Corollary 5.9).
[…]
Remark 1.3.
Most part of this work (except for (2) of Theorem 1.2) was done when the author was a postdoc at Einstein Institute of Mathematics (2017-2018), and versions of that have been put on Arxiv in early of 2019 (see [Xu19b]). Sometime after that, we were aware of that Abdellatif and Hauseux have announced their results on the same problem in which they work for groups of semi-simple rank one ([Abd21]). As far as we know (as of April/2024), their work has not appeared yet, and indeed some of their argument is close to ours.
This paper is organized as follows. In section 2, after setting up general notations, we recall some preliminaries on weights and the Hecke operator $T$. In section 3, we define certain $I_{1,K}$-invariant maps $S_{K}$ and $S_{-}$, and verify their basic properties. In section 4, we prove Theorem 1.1. In section 5, we prove Theorem 1.2.
2 Notations and Preliminaries
2.1 General notations
Let $E/F$ be a unramified quadratic extension of non-archimedean local fields of odd residue characteristic $p$. Let $\mathfrak{o}_{E}$ be the ring of integers of $E$, $\mathfrak{p}_{E}$ be the maximal ideal of $\mathfrak{o}_{E}$, and $k_{E}=\mathfrak{o}_{E}/\mathfrak{p}_{E}$ be the residue field. Fix a uniformizer $\varpi_{E}$ in $E$. Equip $E^{3}$ with the Hermitian form h:
$\text{h}:~{}E^{3}\times E^{3}\rightarrow E$, $(v_{1},v_{2})\mapsto~{}v_{1}^{\text{T}}\beta\overline{v_{2}},v_{1},v_{2}\in E^{3}$.
Here, -- is a generator of $\text{Gal}(E/F)$, and $\beta$ is the matrix
$\begin{matrix}\begin{pmatrix}0&0&1\\ 0&1&0\\ 1&0&0\end{pmatrix}\end{matrix}.$
The unitary group $G$ is defined as:
$G=\{g\in\text{GL}(3,E)\mid\text{h}(gv_{1},gv_{2})=\text{h}(v_{1},v_{2}),\forall~{}v_{1},v_{2}~{}\in E^{3}\}.$
Let $B=HN$ (resp, $B^{\prime}=HN^{\prime}$) be the subgroup of upper (resp, lower) triangular matrices of $G$, with $N$ (resp, $N^{\prime}$) the unipotent radical of $B$ (resp, $B^{\prime}$) and $H$ the diagonal subgroup of $G$. A typical element in $H$ is of the following form and is denoted by $h(x,y)$:
$h(x,y)=\begin{pmatrix}x&0&0\\ 0&y&0\\ 0&0&\bar{x}^{-1}\end{pmatrix}$
for $x\in E^{\times},y\in E^{1}$. We will write $h(x,-\bar{x}x^{-1})$ as $h(x)$ for short. An element in $N$ and $N^{\prime}$ is of the following form
$\begin{pmatrix}1&x&y\\ 0&1&-\bar{x}\\ 0&0&1\end{pmatrix}$,   $\begin{pmatrix}1&0&0\\ x&1&0\\ y&-\bar{x}&1\end{pmatrix}$
and are denoted by $n(x,y)$ and $n^{\prime}(x,y)$. Here we recall that $(x,y)\in E^{2}$ satisfies the relation $x\bar{x}+y+\bar{y}=0$.
For any $k\in\mathbb{Z}$, denote by $N_{k}$ and $N^{\prime}_{k}$ respectively the following subgroups of $N$ and $N^{\prime}$
$N_{k}=\{n(x,y)\in N\mid y\in\mathfrak{p}^{k}_{E}\}$,
$N^{\prime}_{k}=\{n^{\prime}(x,y)\in N^{\prime}\mid y\in\mathfrak{p}^{k}_{E}\}$.
We record a useful identity in $G$: for $y\neq 0$,
$\beta n(x,y)=n(\bar{y}^{-1}x,y^{-1})\cdot h(\bar{y}^{-1})\cdot n^{\prime}(-\bar{y}^{-1}\bar{x},y^{-1}).$ (1)
Up to conjugacy, the group $G$ has two maximal compact open subgroups $K_{0}$ and $K_{1}$, given by:
$K_{0}=\begin{pmatrix}\mathfrak{o}_{E}&\mathfrak{o}_{E}&\mathfrak{o}_{E}\\ \mathfrak{o}_{E}&\mathfrak{o}_{E}&\mathfrak{o}_{E}\\ \mathfrak{o}_{E}&\mathfrak{o}_{E}&\mathfrak{o}_{E}\end{pmatrix}\cap G,~{}K_{1}=\begin{pmatrix}\mathfrak{o}_{E}&\mathfrak{o}_{E}&\mathfrak{p}^{-1}_{E}\\ \mathfrak{p}_{E}&\mathfrak{o}_{E}&\mathfrak{o}_{E}\\ \mathfrak{p}_{E}&\mathfrak{p}_{E}&\mathfrak{o}_{E}\end{pmatrix}\cap G.$
The maximal normal pro-$p$ subgroups of $K_{0}$ and $K_{1}$ are respectively:
$K^{1}_{0}=1+\varpi_{E}M_{3}(\mathfrak{o}_{E})\cap G,~{}K^{1}_{1}=\begin{pmatrix}1+\mathfrak{p}_{E}&\mathfrak{o}_{E}&\mathfrak{o}_{E}\\ \mathfrak{p}_{E}&1+\mathfrak{p}_{E}&\mathfrak{o}_{E}\\ \mathfrak{p}^{2}_{E}&\mathfrak{p}_{E}&1+\mathfrak{p}_{E}\end{pmatrix}\cap G.$
Let $\alpha$ be the following diagonal matrix in $G$:
$\begin{matrix}\begin{pmatrix}\varpi_{E}^{-1}&0&0\\ 0&1&0\\ 0&0&\varpi_{E}\end{pmatrix}\end{matrix},$
and put $\beta^{\prime}=\beta\alpha^{-1}$. Note that $\beta\in K_{0}$ and $\beta^{\prime}\in K_{1}$. We use $\beta_{K}$ to denote the unique element in $K\cap\{\beta,\beta^{\prime}\}$.
Let $K\in\{K_{0},K_{1}\}$, and $K^{1}$ be the maximal normal pro-$p$ subgroup of $K$. We identify the finite group $\Gamma_{K}=K/K^{1}$ with the $k_{F}$-points of an algebraic group defined over $k_{F}$. Let $\mathbb{B}$ (resp, $\mathbb{B}^{\prime}$) be the upper (resp, lower) triangular subgroup of $\Gamma_{K}$, and $\mathbb{U}$ (resp, $\mathbb{U}^{\prime}$) be its unipotent radical. The Iwahori subgroup $I_{K}$ (resp, $I^{\prime}_{K}$) and pro-$p$ Iwahori subgroup $I_{1,K}$ (resp, $I^{\prime}_{1,K}$) in $K$ are the inverse images of $\mathbb{B}$ (resp, $\mathbb{B}^{\prime}$) and $\mathbb{U}$ (resp, $\mathbb{U}^{\prime}$).
Denote by $n_{K}$ and $m_{K}$ the unique integers such that $N\cap I_{1,K}=N_{n_{K}}$ and $N^{\prime}\cap I_{1,K}=N^{\prime}_{m_{K}}$. We have $n_{K}+m_{K}=1$. Note that the coset spaces $N_{n_{K}}/N_{n_{K}+1}$ and $N^{\prime}_{m_{K}}/N^{\prime}_{m_{K}+1}$ are indeed groups of order respectively $q^{t_{K}}$ and $q^{4-t_{K}}$, where $t_{K}=3$ or $1$, depending on $K$ is hyperspecial or not.
All representations in this note are smooth over $\overline{\mathbf{F}}_{p}$.
2.2 Weights
Let $\sigma$ be an irreducible smooth representation of $K$. As $K_{1}$ is pro-$p$ and normal in $K$, $\sigma$ factors through the finite group $\Gamma_{K}$, i.e., $\sigma$ is the inflation of an irreducible representation of $\Gamma_{K}$. Conversely, any irreducible representation of $\Gamma_{K}$ inflates to an irreducible smooth representation of $K$. We may therefore identify irreducible smooth representations of $K$ with irreducible representations of $\Gamma_{K}$, and we shall call them weights of $K$ or $\Gamma_{K}$ from now on.
For a weight $\sigma$ of $K$, it is known that $\sigma^{I_{1,K}}$ and $\sigma_{I^{\prime}_{1,K}}$ are one-dimensional, and that the natural composition map $\sigma^{I_{1,K}}\hookrightarrow\sigma\twoheadrightarrow\sigma_{I^{\prime}_{1,K}}$ is an isomorphism of vector spaces ([CE04, Theorem 6.12]). This implies there exists a unique $\lambda_{\beta_{K},\sigma}\in\overline{\mathbf{F}}_{p}$, such that $\beta_{K}\cdot v-\lambda_{\beta_{K},\sigma}v\in\sigma(I^{\prime}_{1,K})$, for $v\in\sigma^{I_{1,K}}$. By [HV15, Proposition 3.16] and the fact that $\beta_{K}\notin I_{K}\cdot I^{\prime}_{K}$, the scalar $\lambda_{\beta_{K},\sigma}$ is zero if $\text{dim}\sigma>1$, and is equal to $\sigma(\beta_{K})$ if $\text{dim}\sigma=1$.
2.3 The Hecke operator $T$
Let $K\in\{K_{0},K_{1}\}$, and $\sigma$ be a weight of $K$. Let $\text{ind}_{K}^{G}\sigma$ be the maximal compact induction and $\mathcal{H}(K,\sigma):=\text{End}_{G}(\text{ind}_{K}^{G}\sigma)$ be the associate spherical Hecke algebra. The algebra $\mathcal{H}(K,\sigma)$ is isomorphic to $\overline{\mathbf{F}}_{p}[T]$, for certain $T\in\mathcal{H}(K,\sigma)$ ([Her11, Corollary 1.3], see also [Xu19a, Proposition 3.3]).
We don’t recall the exact definition of $T$ but only its formula on a specific function. For a non-zero vector $v\in\sigma$, denote by $\hat{f}_{v}$ the function in $\text{ind}^{G}_{K}\sigma$ supported on $K$ and having value $v$ at $Id$.
Proposition 2.1.
Take a non-zero vector $v_{0}$ in $\sigma^{I_{1,K}}$. Then, we have
$T\hat{f}_{v_{0}}=\sum_{u\in N_{n_{K}}/N_{n_{K}+2}}~{}u\alpha^{-1}\cdot\hat{f}_{v_{0}}+\lambda_{\beta_{K},\sigma}\sum_{u\in N_{n_{K}+1}/N_{n_{K}+2}}\beta_{K}u\alpha^{-1}\cdot\hat{f}_{v_{0}}$ (2)
Proof.
This is [Xu19a, Proposition 3.6]. Note that the above formula determines $T$ uniquely, as the function $\hat{f}_{v_{0}}$ generates the whole representation $\text{ind}^{G}_{K}\sigma$. ∎
3 The $I_{1,K}$-invariant maps $S_{K}$ and $S_{-}$
In this section, we recall some partial linear operators on a smooth representation $\pi$, and their certain invariant properties.
Definition 3.1.
Let $\pi$ be a smooth representation of $G$. We define:
$S_{K}:\pi^{N^{\prime}_{m_{K}}}\rightarrow\pi^{N_{n_{K}}}$,
$v\mapsto\sum_{u\in N_{n_{K}}/N_{n_{K}+1}}~{}u\beta_{K}v$.
$S_{-}:\pi^{N_{n_{K}}}\rightarrow\pi^{N^{\prime}_{m_{K}}}$,
$v\mapsto\sum_{u\in N_{n_{K}+1}/N_{n_{K}+2}}~{}\beta_{K}u\alpha^{-1}v$
It is simple to check both $S_{K}$ and $S_{-}$ are well-defined.
Proposition 3.2.
We have:
$(1)$.  Let $h\in H_{1}=I_{1,K}\cap H$. Then $S_{K}(hv)=h^{s}\cdot S_{K}v$, for $v\in\pi^{N^{\prime}_{m_{K}}}$, and $S_{-}(hv)=h^{s}\cdot S_{-}v$, , for $v\in\pi^{N_{n_{K}}}$, where $h^{s}$ is short for $\beta_{K}h\beta_{K}$.
$(2)$.  If $v$ is fixed by $I_{1,K}$, the same is true for $S_{K}\cdot v$ and $S_{-}\cdot v$.
Proof.
This is firstly proved in [Xu17]. We reproduce the arguemnt in full here.
For $(1)$, we note that the group $H_{1}$ acts on $\pi^{N_{n_{K}}}$ and $\pi^{N^{\prime}_{m_{K}}}$, as it normalizes $N_{n_{K}}$ and $N^{\prime}_{m_{K}}$. The statement then follows from the definitions.
To prove $(2)$, we need the following Lemma.
Lemma 3.3.
Given a $u^{\prime}\in N^{\prime}_{m_{K}}$ and a $u\in N_{n_{K}}$.
$(1)$. There is a unique $u_{1}\in N_{n_{K}},h\in H_{1},u^{\prime}_{1}\in N^{\prime}_{m_{K}}$ so that the following identity
$u^{\prime}u=u_{1}hu^{\prime}_{1}$
holds.
$(2)$. For any $l>m\geq 0$, when $u$ goes through $N_{n_{K}+m}/N_{n_{K}+l}$, the element $u_{1}$ also goes through $N_{n_{K}+m}/N_{n_{K}+l}$.
Proof.
The uniqueness statement is clear, and only the existence needs to be proved. Assume $u=n(x,y)\in N,u^{\prime}\in n^{\prime}(x^{\prime},y^{\prime})\in N^{\prime}$. Then, if $1+xx^{\prime}+\overline{yy^{\prime}}\in E^{\times}$, we have
$u^{\prime}u=u_{1}hu^{\prime}_{1}$,
where $u_{1}=n(x_{1},y_{1})\in N$ in which $x_{1},y_{1}$ are given by
$x_{1}=\frac{x-\overline{yx^{\prime}}}{1+xx^{\prime}+\overline{yy^{\prime}}},y_{1}=\frac{y}{1+\overline{xx^{\prime}}+yy^{\prime}}$,
and $h$ is the diagonal matrix
$\begin{pmatrix}\frac{1}{1+xx^{\prime}+\overline{yy^{\prime}}}&0&0\\ 0&\frac{1+xx^{\prime}+\overline{yy^{\prime}}}{1+\overline{xx^{\prime}}+yy^{\prime}}&0\\ 0&0&1+\overline{xx^{\prime}}+yy^{\prime}\end{pmatrix},$
and $u^{\prime}_{1}=n^{\prime}(x^{\prime}_{1},y^{\prime}_{1})\in N^{\prime}$ in which
$x^{\prime}_{1}=\frac{x^{\prime}-\overline{xy^{\prime}}}{1+xx^{\prime}+\overline{yy^{\prime}}},y^{\prime}_{1}=\frac{y^{\prime}}{1+\overline{xx^{\prime}}+yy^{\prime}}$.
Under our assumption that $u^{\prime}\in N^{\prime}_{m_{K}}$ and $u\in N_{n_{K}}$, the condition $1+xx^{\prime}+\overline{yy^{\prime}}\in E^{\times}$ holds automatically. The existence is established.
We continue to prove $(2)$. From the formulae of $x_{1}$ and $y_{1}$ given above, one checks by a direct computation that
$u_{1}=uN_{n_{K}+m+1}$, if $u\in N_{n_{K}+m}$ for some $m\geq 0$.
Explicitly,
$u_{1}=u\cdot n(\ast,yz)$ (3)
holds for some $z\in\mathfrak{p}_{E}$. Recall that $u=n(x,y)\in N_{n_{K}},u^{\prime}=n^{\prime}(x^{\prime},y^{\prime})\in N^{\prime}_{m_{K}}$.
We may therefore view $u^{\prime}$ as a map
$u^{\prime}:N_{n_{K}+m}/N_{n_{K}+l}\rightarrow N_{n_{K}+m}/N_{n_{K}+l}$
$uN_{n_{K}+l}\mapsto u_{1}N_{n_{K}+l}$
It suffices to show the map is injective. Assume for another $w\in N_{n_{K}}$, we have a decomposition $u^{\prime}w=u_{2}b^{\prime\prime}$ for $u_{2}\in N_{n_{K}}$ and $b^{\prime\prime}\in B^{\prime}$. We have to prove:
$u_{2}\in u_{1}N_{n_{K}+l}$ implies $w\in uN_{n_{K}+l}$.
Write $u^{-1}_{1}u_{2}$ as $u_{3}$. A little algebraic transform gives:
$w=u\cdot b^{\prime-1}u_{3}b^{\prime\prime}$
We need to check that the element $b^{\prime-1}u_{3}b^{\prime\prime}\in N_{n_{K}}$, denoted by $u_{4}$, lies in $N_{n_{K}+l}$. The element $b^{\prime}$ can be written as $h\cdot u^{\prime}_{1}$, for a diagonal matrix $h\in H_{1}$ and $u^{\prime}_{1}\in N^{\prime}_{m_{K}}$. We therefore get
$u^{\prime}_{1}u_{4}=(h^{-1}u_{3}h)\cdot h^{-1}b^{\prime\prime}$,
where the right hand side is a decomposition of $u^{\prime}_{1}u_{4}$ given in $(1)$. The uniqueness of such a decomposition implies
$u_{4}\in N_{n_{K}+l}$ if and only if $h^{-1}u_{3}h\in N_{n_{K}+l}$
for any $l\geq 0$. Our assumption is that $u_{3}=u^{-1}_{1}u_{2}\in N_{n_{K}+l}$, which is the same as $h^{-1}u_{3}h\in N_{n_{K}+l}$ ($h\in H_{1}$). We are done. ∎
We proceed to complete the argument of $(2)$ of the Proposition. By $(1)$ and the decomposition of $I_{1,K}=N^{\prime}_{m_{K}}\times H_{1}\times N_{n_{K}}$, it suffices to check that, for $u^{\prime}=n^{\prime}(x,y)\in N^{\prime}_{m_{K}}$, the element $u^{\prime}\cdot S_{K}v$
$u^{\prime}\cdot S_{K}v=\sum_{u\in N_{n_{K}}/N_{n_{K}+1}}~{}u^{\prime}u\beta_{K}v$
is still equal to $S_{K}v=\sum_{u\in N_{n_{K}}/N_{n_{K}+1}}~{}u\beta_{K}v$. By $(1)$ of Lemma 3.3, the right hand side of above sum is equal to:
$\sum_{u\in N_{n_{K}}/N_{n_{K}+1}}~{}u_{1}hu^{\prime}_{1}\beta_{K}v$.
We get:
$u^{\prime}\cdot S_{K}v=\sum_{u\in N_{n_{K}}/N_{n_{K}+1}}~{}u_{1}\beta_{K}(\beta_{K}hu^{\prime}_{1}\beta_{K})v=\sum_{u\in N_{n_{K}}/N_{n_{K}+1}}~{}u_{1}\beta_{K}v$,
which, by $(2)$ of Lemma 3.3, is just $\sum_{u_{1}\in N_{n_{K}}/N_{n_{K}+1}}~{}u_{1}\beta_{K}v$. The argument for the statement $S_{K}v\in\pi^{I_{1,K}}$ for $v\in\pi^{I_{1,K}}$ is now complete.
By almost the same argument, one can verify that $S_{-}\cdot v\in\pi^{I_{1,K}}$ for $v\in\pi^{I_{1,K}}$. ∎
Remark 3.4.
A slight variant of $(2)$ holds by the same argument. When $u$ goes through $N_{n_{K}+m}\setminus N_{n_{K}+n}/N_{n_{K}+l}$, the element $u_{1}$ also goes through $N_{n_{K}+m}\setminus N_{n_{K}+n}/N_{n_{K}+l}$, for any $l>n\geq m\geq 0$.
We apply the operators $S_{-}$ and $S_{K}$ to the $I_{1,K}$-invariants of a principal series $\text{ind}^{G}_{B}\varepsilon$. Recall the space $(\text{ind}^{G}_{B}\varepsilon)^{I_{1,K}}$ is spanned by the functions $g_{1}$ and $g_{2}$ which are characterized by: $g_{1}(Id)=1,g_{1}(\beta_{K})=0,g_{2}(Id)=0,g_{2}(\beta_{K})=1$.
Proposition 3.5.
We have:
$(1)$$S_{K}g_{1}=g_{2},~{}S_{K}g_{2}=d_{q^{t_{K}}}\cdot g_{2}$.
$(2)$$S_{-}g_{1}=d_{q^{4-t_{K}}}\cdot g_{1},~{}S_{-}g_{2}=\varepsilon(\alpha)g_{1}$.
Here, $d_{q^{t_{K}}}=\sum_{(x,t)\in L^{\times}_{q^{t_{K}}}}\varepsilon_{0}((h(t)),d_{q^{4-t_{K}}}=\sum_{(x,t)\in L^{\times}_{q^{4-t_{K}}}}\varepsilon_{0}((h(t))$, $\varepsilon_{0}=\varepsilon\mid_{B\cap K}$.
Proof.
By Proposition 3.2, it suffices to compute the values of the functions in consideration at $Id$ and $\beta_{K}$. We omit the details. Note that the exact values of $d_{q^{t_{K}}}$ and $d_{q^{4-t_{K}}}$ depend on the nature of $\varepsilon_{0}$, see [KX15, Appendix A]. ∎
Later on, we will use the composition $S_{+}=S_{K}\circ S_{-}$, which lies in $\text{End}_{\overline{\mathbf{F}}_{p}}(\pi^{N_{n_{K}}})$. Explicitly for $v\in\pi^{N_{n_{K}}}$, we have
$S_{+}v=\sum_{u\in N_{n_{K}}/N_{n_{K}+2}}~{}u\alpha^{-1}v$.
By Proposition 3.2, it preserves the $I_{1,K}$-invariants of a smooth representation.
4 Non-supersingular representations
4.1 Some recaps
In this subsection, we recall briefly the restriction of a principal series to a Borel subgroup. For readers’ convenience, we reproduce certain details below where we mainly follow the approach in [Vig08].
For a character $\varepsilon$ of $B$, consider the principal series $\text{ind}^{G}_{B}\varepsilon$. Recall that $\text{ind}^{G}_{B}\varepsilon$ is reducible if and only if $\varepsilon=\eta\circ\text{det}$ for some character $\eta$, and in this case it is of length two. Evaluating an $f\in\text{ind}^{G}_{B}\varepsilon$ at the identity, we get a $B$-map from the principal series to the character $\varepsilon$. Denote the kernel by $\kappa_{\varepsilon}$. Then we have a short exact sequence of $B$-representations:
$0\rightarrow\kappa_{\varepsilon}\rightarrow\text{ind}^{G}_{B}\varepsilon\rightarrow\varepsilon\rightarrow 0$
By almost the same argument of [Vig08, Theorem 5], one shows that the $B$-representation $\kappa_{\varepsilon}$ is irreducible. Indeed, as explained below, one may even prove $\kappa_{\varepsilon}$ is irreducible as a representation of $\alpha^{\mathbb{Z}}N$. But by the same map $\Phi$ below, one may verify that $St\mid_{B}\cong\kappa_{1}$. This gives irreducibility of $St\mid_{B}$.
Lemma 4.1.
The $B$-representation $\kappa_{\varepsilon}$ is irreducible.
Proof.
1).  We firstly identify the underlying space of $\kappa_{\varepsilon}$ with $C^{\infty}_{c}(N)$.
$\Phi:\kappa_{\varepsilon}\rightarrow C^{\infty}_{c}(N),f\mapsto\Phi(f),\Phi(f)(u)=f(\beta u),\forall u\in N.$
$\Psi:C^{\infty}_{c}(N)\rightarrow\kappa_{\varepsilon},f\mapsto\Psi(f),\Psi(f)(b\beta u)=\varepsilon(b)f(u),\forall b\in B,u\in N.$
One verifies easily that $\Psi$ and $\Phi$ are inverse to each other. This gives $C^{\infty}_{c}(N)$ a structure of $B$-representation.
2). We modify the argument of [Ly15, Proposition 5.2] to our case. Let $V$ be a non-zero $B$-stable subspace of $C^{\infty}_{c}(N)$, and $f$ be a non-zero function in $V$. As $f$ is compactly supported and $N$ has the decreasing open compact cover $(N_{k})_{k\in\mathbb{Z}}$, we may assume the support of $f$ is contained in $N_{k}$ for some integer $k$. Write $V_{k}$ for the subspace of $V$ consisting of functions supported in $N_{k}$, we have $V_{k}\neq 0$. By [BL95, Lemma 1], we know $V^{N_{k}}_{k}\neq 0$. This shows that $V$ contains the characteristic function $1_{N_{k}}$ of $N_{k}$.
Now for any $n\in\mathbb{Z},u\in N$, we have $u\alpha^{n}\cdot 1_{N_{k}}=\varepsilon(\alpha^{-n})1_{N_{k-2n}u^{-1}}$, and as $V$ is $B$-stable we conclude $V$ contains $1_{N_{k-2n}u^{-1}}$. Note that $1_{N_{k-1}}=\sum_{u\in N_{k-1}/N_{k}}u^{-1}\cdot 1_{N_{k}}$, so we have $1_{N_{k-1}}\in V$. Repeating the previous process, we conclude $V$ contains $1_{N_{k-2n-1}u^{-1}}$. In all we have shown $V$ contains all the functions $1_{N_{k}u}$ for any $k\in\mathbb{Z}$ and $u\in N$. As all the functions $\{1_{N_{k}u}\mid k\in\mathbb{Z},u\in N\}$ span the underlying space of $C^{\infty}_{c}(N)$, we get $V=C^{\infty}_{c}(N)$. ∎
4.2 Proof of Theorem 1.1
We now come to the main input of this section.
Theorem 4.2.
The restriction map induces an isomorphism between the following spaces:
$\textnormal{Hom}_{G}(\textnormal{ind}^{G}_{B}\varepsilon,\pi)\cong\textnormal{Hom}_{B}(\kappa_{\varepsilon},\pi)$
Proof.
We show firstly that the restriction map is injective.
Given $\phi_{1}$ and $\phi_{2}$ in the space $\textnormal{Hom}_{G}(\textnormal{ind}^{G}_{B}\varepsilon,\pi)$, suppose that $\phi=\phi_{1}-\phi_{2}$ vanishes at the subspace $\kappa_{\varepsilon}$. By the remark proceeding Lemma 4.1, $\phi$ induces a $B$-map from the character $\varepsilon$ to $\pi$, for which we still denote by $\phi$.
Lemma 4.3.
If $\varepsilon\neq\eta\circ\textnormal{det}$ for any character $\eta$, then $\textnormal{Hom}_{B}(\varepsilon,\pi)=0$.
Proof.
Assume $\phi\neq 0$. As $\pi$ is smooth, the vector $\phi(1)\in\pi$ is fixed by some $N^{\prime}_{m_{K}+2k}$ for large enough $k$. Using the following identity repeatedly
$\alpha N^{\prime}_{m_{K}+2k-2}\alpha^{-1}=N^{\prime}_{m_{K}+2k}$
and $\phi(\alpha\cdot 1)=\varepsilon(\alpha)\phi(1)=\alpha\cdot\phi(1)$, we see $\phi(1)$ is fixed by $N^{\prime}$. As the group $G$ is generated by $B$ and $N^{\prime}$, we see $\varepsilon$ extends uniquely to a character of $G$ (put $\varepsilon(N^{\prime})=1$). In such a situation, $\textnormal{Hom}_{B}(\varepsilon,\pi)\cong\textnormal{Hom}_{G}(\varepsilon,\pi)$. ∎
Remark 4.4.
Under the same assumption on $\varepsilon$, the Lemma implies that any non-zero map in $\textnormal{Hom}_{B}(\textnormal{ind}^{G}_{B}\varepsilon,\pi)$ is an injection. Take $\pi=\textnormal{ind}^{G}_{B}\varepsilon$. We conclude that $\textnormal{End}_{B}(\textnormal{ind}^{G}_{B}\varepsilon)$ is one-dimensional. This is because, by the proceeding remark, any non-zero map in the former space will induce a non-zero map in $\textnormal{End}_{B}(\kappa_{\varepsilon})$ which is one-dimensional by Lemma 4.1. We deduce that $\textnormal{End}_{B}(\textnormal{ind}^{G}_{B}\varepsilon)\cong\textnormal{End}_{G}(\textnormal{ind}^{G}_{B}\varepsilon)$.
We are done if $\varepsilon\neq\eta\circ\text{det}$ for any character $\eta$. Assume $\varepsilon=\eta\circ\text{det}$ for some $\eta$. After a twist we may assume $\eta=1$. If $\phi\neq 0$, it induces a non-zero map in $\textnormal{Hom}_{B}(1,\pi)\cong\textnormal{Hom}_{G}(1,\pi)$. This in turn implies the map $\phi\in\textnormal{Hom}_{G}(\textnormal{ind}^{G}_{B}1,\pi)$ realizes the trivial character of $G$ as a quotient of $\textnormal{ind}^{G}_{B}1$ (note that $\phi$ is not injective by our assumption), which is not true.
We proceed to prove the restriction map is surjective.
Recall again that the space $(\textnormal{ind}^{G}_{B}\varepsilon)^{I_{1,K}}$ is two dimensional with a basis of functions $g_{1}$ and $g_{2}$ characterized by: $g_{1}(Id)=1,g_{1}(\beta_{K})=0,g_{2}(Id)=0,g_{2}(\beta_{K})=1$. By Proposition 3.5, we have
$S_{+}g_{2}=\varepsilon(\alpha)g_{2}$ (4)
Then, by Lemma 5.3 the $K$-representation $\langle K\cdot g_{2}\rangle=\langle K\cdot S_{+}g_{2}\rangle$ is a weight, denoted by $\sigma$, of dimension greater than one (note that $I_{K}$ acts on $g_{2}$ by the character $\varepsilon^{s}_{0}$).
Let $\phi$ be a non-zero $B$-map from $\kappa_{\varepsilon}$ to $\pi$. The function $g_{2}$ by definition is supported on $B\beta_{K}I_{K}$ so it lies in $\kappa_{\varepsilon}$. As $\kappa_{\varepsilon}$ is irreducible (Lemma 4.1), we have $\phi(g_{2})$ is non-zero. Since $\phi$ respects the action of $B$, the vector $\phi(g_{2})$ is fixed by $B\cap I_{1,K}$. Now we compute $\phi(S_{+}g_{2})$:
$\phi(S_{+}g_{2})=\varepsilon(\alpha)\phi(g_{2})=S_{+}\phi(g_{2})$
that is
$\phi(g_{2})=\varepsilon(\alpha)^{-1}S_{+}\phi(g_{2})$ (5)
As $\pi$ is smooth, the vector $\phi(g_{2})$ is fixed by some $N^{\prime}_{m_{K}+2k}$ for $k$ large enough. Now by applying the argument of Lemma 3.3 to the above equality, we see $\phi(g_{2})$ is fixed by $N^{\prime}_{m_{K}+2k-2}$. Repeating such a process enough times, we prove that the vector $\phi(g_{2})$ is fixed by $N^{\prime}_{m_{K}}$. By Iwahori decomposition $I_{1,K}=N^{\prime}_{m_{K}}\cdot(B\cap I_{1,K})$, we conclude that $\phi(g_{2})$ is fixed by $I_{1,K}$.
By Lemma 5.3 again the representation $\langle K\cdot\phi(g_{2})\rangle$ is a weight $\sigma^{\prime}$ of dimension greater than one. We claim that $\sigma\cong\sigma^{\prime}$. The Iwahori group $I_{K}$ acts on the vector $\phi(g_{2})$ by $\varepsilon^{s}_{0}$. If $\varepsilon_{0}\neq\eta\circ\text{det}$ for $K=K_{0}$ (resp, $\varepsilon_{0}\neq\varepsilon^{s}_{0}$ for $K=K_{1}$), the claim follows as in this case a weight is determined by the character of $I_{K}$ acting on its $I_{1,K}$-invariants. If it is in the other case, we are also done: neither $\sigma$ or $\sigma^{\prime}$ is a one-dimensional character, and as quotients of the principal series $\text{Ind}^{K}_{I_{K}}\varepsilon^{s}_{0}$ they are both isomorphic to $st\otimes\varepsilon_{0}$.
By [Xu18b, Proposition 4.14, 4.15], we have an isomorphism, unique up to a scalar,
$\text{ind}^{G}_{K}\sigma/(T_{\sigma}-\varepsilon(\alpha))\cong\text{ind}^{G}_{B}\varepsilon$.
As the representation $\langle G\cdot\phi(g_{2})\rangle$ contains the weight $\sigma$, it is therefore a quotient of the above representation (by (5) and that $\sigma$ is not a character). We have shown $\phi$ extends to a $G$-map as required. ∎
With the last Theorem proved, we conclude with the following corollaries.
Corollary 4.5.
Suppose $\varepsilon\neq\eta\circ\textnormal{det}$ for any character $\eta$. Then
$\textnormal{Hom}_{G}(\textnormal{ind}^{G}_{B}\varepsilon,\pi)\cong\textnormal{Hom}_{B}(\textnormal{ind}^{G}_{B}\varepsilon,\pi)$
Proof.
Let $\phi\in\textnormal{Hom}_{B}(\textnormal{ind}^{G}_{B}\varepsilon,\pi)$ be non-zero. By Remark 4.4, we know $\phi$ is injective. We have:
1).  Using Remark 4.4 again, the image of $\phi$ is contained in $\langle G\cdot\phi(\kappa_{\varepsilon})\rangle$.
2).  By Theorem 4.2 and irreducibility of $\textnormal{ind}^{G}_{B}\varepsilon$, there is an isomorphism $\text{ind}^{G}_{B}\varepsilon\cong\langle G\cdot\phi(\kappa_{\varepsilon})\rangle$.
With these ingredients available, the proof is then identical to that of [Paš07, Corollary 5.5]. ∎
Corollary 4.6.
We have
$\textnormal{Hom}_{B}(St,\pi)\cong\textnormal{Hom}_{G}(\textnormal{ind}^{G}_{B}1,\pi)$
Proof.
As $St\mid_{B}\cong\kappa_{1}$ (remarks before Lemma 4.1), the assertion is a special case of Theorem 4.2. Note that the result can not be improved by replacing $\textnormal{ind}^{G}_{B}1$ in the statement by $St$: the space $\textnormal{Hom}_{G}(\textnormal{ind}^{G}_{B}1,\textnormal{ind}^{G}_{B}1)\neq 0$ but $\textnormal{Hom}_{G}(St,\textnormal{ind}^{G}_{B}1)=0$. ∎
5 Supersingular representations
5.1 Definition
Recall we have defined the Hecke operator $T$ in subsection 2.3. To define the supersingular representations, we modify it in the following way. If $\textnormal{dim}~{}\sigma=1$, we put $T_{\sigma}=T+1$; otherwise, we put $T_{\sigma}=T$. Note that this modification reflects the fact the group $G$ has two maximal compact open subgroups, up to conjugacy.
Definition 5.1.
An irreducible smooth $\overline{\mathbf{F}}_{p}$-representation $\pi$ of $G$ is called supersingular if it is a quotient of $\textnormal{ind}^{G}_{K}\sigma/(T_{\sigma})$, for some weight $\sigma$ of $K$.
5.2 A key property
Let $\pi$ be an irreducible smooth representation of $G$, and $\sigma$ be a weight of $K$ contained in $\pi$. By [Xu18a, Theorem 1.1], the representation $\pi$ admits Hecke eigenvalues for the spherical Hecke algebra $\mathcal{H}(K,\sigma)\cong\overline{\mathbf{F}}_{p}[T_{\sigma}]$. This implies that the representation $\pi$ is a quotient of $\text{ind}^{G}_{K}\sigma/(T_{\sigma}-\lambda)$, for some scalar $\lambda$. By [Xu18b, Theorem 1.1], the representation $\pi$ is supersingular if $\lambda=0$.
Lemma 5.2.
Let $\pi$ be a supersingular representation of $G$, and assume $\phi$ is a non-zero $G$-map from $\textnormal{ind}^{G}_{K}\sigma$ to $\pi$. Then, for large enough $k\geq 1$, we have
$\phi\circ T_{\sigma}^{k}=0.$
Proof.
By [Xu18a, Corollary 4.2], there is a non-constant polynomial $P(X)$ such that $\phi\circ P(T_{\sigma})=0$. Assume $P(X)$ is such a polynomial of minimal degree. Take a root $\lambda$ of $P(X)$, and write $P(X)=(X-\lambda)P_{1}(X)$. Put $\phi^{\prime}=\phi\circ P_{1}(T_{\sigma})$. Note that $\phi^{\prime}$ is still a $G$-map from $\textnormal{ind}^{G}_{K}\sigma$ to $\pi$. By our assumption, the map $\phi^{\prime}$ is non-zero and factors through $\textnormal{ind}^{G}_{K}\sigma/(T_{\sigma}-\lambda)$. As $\pi$ is supersingular, we have $\lambda=0$ ([Xu18b, Theorem 1.1]). We conclude $P(X)=X^{n}$ for some $n\geq 1$. ∎
Lemma 5.3.
Let $\pi$ be a smooth representation of $G$. Assume $v$ is a non-zero vector in $\pi^{I_{1,K}}$, such that $I_{K}$ acts on $v$ as a character. Then, either $S_{+}v=0$, or $S_{+}v$ generates a weight of $K$ of dimension greater than one.
Proof.
Assume $S_{+}v\neq 0$. We put $w=S_{-}v$. By definition,
$S_{+}v=S_{K}w$,
and we see $w$ must be non-zero. Consider the $K$-representation $\kappa=\langle K\cdot w\rangle$. As $I_{K}$ acts on $v$ by a character $\chi$, $I_{K}$ acts on $w$ by $\chi^{s}$ (Proposition 3.2). By Frobenius reciprocity, there is a surjective $K$-map from $\text{Ind}^{K}_{I_{K}}\chi^{s}$ to $\kappa$, sending $\varphi_{\chi^{s}}$ to $w$. Here, $\varphi_{\chi^{s}}$ is the function in $\text{Ind}^{K}_{I_{K}}\chi^{s}$ supported on $I_{K}$ and having value $1$ at $Id$.
Via aforementioned map, we see $\langle K\cdot S_{+}v\rangle$ is the image of $\langle K\cdot S_{K}\varphi_{\chi^{s}}\rangle$. But the latter, by [KX15, Proposition 5.7], is an irreducible smooth representation of $K$ of dimension greater than one. The assertion follows. ∎
Proposition 5.4.
Assume $\pi$ is a supersingular representation of $G$, and $v$ is a non-zero vector in $\pi^{I_{1,K}}$. Then, for $k\gg 0$, we have $S^{k}_{+}v=0$.
Proof.
Assume firstly $I_{K}$ acts on $v$ as a character $\chi$.
Assume $S_{+}v\neq 0$. By Lemma 5.3 the $K$-subrepresentation generated by $S_{+}v$ is a weight of dimension greater than one, and denote it by $\sigma$. By Frobenius reciprocity, we have a $G$-map $\phi$ from $\textnormal{ind}^{G}_{K}\sigma$ to $\pi$, sending the function $\hat{f}_{S_{+}v}$ to $S_{+}v$. From Lemma 5.2, there is some $k\geq 1$ such that
$S_{+}^{k}v=0,$
and we are done in this special case.
Note that $I_{K}/I_{1,K}$ is an abelian group of finite order prime to $p$. For any non-zero $v\in\pi^{I_{1,K}}$, the $I_{K}$-representation $\langle I_{K}\cdot v\rangle$ generated by $v$ is a finite sum of characters, and we may write $v$ as $\sum v_{i}$ so that $I_{K}$ acts on $v_{i}$ by a character $\chi_{i}$ of $I_{K}/I_{1,K}$. We then apply the previous process to each $v_{i}$, and take the largest $k_{i}$. We are done. ∎
[…]
In this part, for an irreducible smooth representation we prove a sufficient condition under which its restriction to the Borel subgroup remains irreducible. We will verify it for supersingular ones in the next part.
Proposition 5.5.
Let $\pi$ be an irreducible smooth representation of $G$. If, for any non-zero vector $w\in\pi$, there is a non-zero vector $v\in\pi^{I_{1,K}}\cap\langle B\cdot w\rangle$ such that
$S_{+}v=0$,
then $\pi\mid_{B}$ is irreducible.
Proof.
Let $w$ be a non-zero vector in $\pi$. As $\pi$ is smooth, there exists a $k\geq 0$ such that $w$ is fixed by $N^{\prime}_{2k+m_{K}}$. Hence, the vector $w_{1}=\alpha^{-k}w$ is fixed by $N^{\prime}_{m_{K}}$. Since $I_{1,K}=(I_{1,K}\cap B)\cdot N^{\prime}_{m_{K}}$, we see
$\langle I_{1,K}\cdot w_{1}\rangle=\langle(I_{1,K}\cap B)\cdot w_{1}\rangle.$
As $I_{1,K}$ is pro-$p$, the space $\langle(I_{1,K}\cap B)\cdot w_{1}\rangle$ has non-zero $I_{1,K}$-invariant ([BL95, Lemma 1]). We conclude that $\pi^{I_{1,K}}\cap\langle B\cdot w\rangle\neq 0$.
Lemma 5.6.
If $S_{+}v=0$, then $\beta_{K}v\in\langle B\cdot v\rangle.$
Proof.
By the assumption $S_{+}v=0$, we get
$v=-\alpha\cdot\sum_{u\in(N_{n_{K}}\setminus N_{n_{K}+2})/N_{n_{K}+2}}u\alpha^{-1}v$
or equivalently
$\beta_{K}v=-\sum_{u\in(N_{n_{K}}\setminus N_{n_{K}+2})/N_{n_{K}+2}}\beta_{K}\alpha u\alpha^{-1}v.$ (6)
Applying (1), we see $\beta_{K}\alpha u\alpha^{-1}\in BN^{\prime}_{m_{K}}$, for any $u\in(N_{n_{K}}\setminus N_{n_{K}+2})/N_{n_{K}+2}$. More precisely, for $u=n(\ast,\varpi^{n_{K}}_{E}t)\in N_{n_{K}}\setminus N_{n_{K}+1}$, we have
$\beta_{K}\alpha u\alpha^{-1}=n(\ast,\varpi^{n_{K}+2}_{E}t^{-1})h(\bar{t}^{-1})\alpha^{-2}n^{\prime}(\ast,\varpi^{m_{K}+1}_{E}t^{-1})$;
for $u=n(\ast,\varpi^{n_{K}+1}_{E}t)\in N_{n_{K}+1}\setminus N_{n_{K}+2}$, we have
$\beta_{K}\alpha u\alpha^{-1}=n(\ast,\varpi^{n_{K}+1}_{E}t^{-1})h(\bar{t}^{-1})\alpha^{-1}n^{\prime}(\ast,\varpi^{m_{K}}_{E}t^{-1}).$
That gives $\beta_{K}\alpha u\alpha^{-1}v\in\langle B\cdot v\rangle$ for all $u\in(N_{n_{K}}\setminus N_{n_{K}+2})/N_{n_{K}+2}$. We conclude $\beta_{K}v\in\langle B\cdot v\rangle$ from the above equality (6). For our later purpose, we record (6) in a more explicit form:
$\beta_{K}v=-\sum_{u}u\alpha^{-2}h(\bar{t}^{-1})v-\sum_{u}u\alpha^{-1}h(\bar{t}^{-1})v,$ (7)
where, in the first sum $u=n(\ast,\varpi^{n_{K}+2}_{E}t)$ goes through $(N_{n_{K}+2}\setminus N_{n_{K}+3})/N_{n_{K}+4}$; in the second sum $u=n(\ast,\varpi^{n_{K}+1}_{E}t)$ goes through $(N_{n_{K}+1}\setminus N_{n_{K}+2})/N_{n_{K}+2}$. ∎
We proceed to complete the proof of Proposition 5.5. Choose $0\neq v\in\pi^{I_{1,K}}\cap\langle B\cdot w\rangle$ such that $S_{+}v=0$. The above Lemma says $\beta_{K}v\in\langle B\cdot v\rangle$. As $\pi$ is irreducible, we have $\pi=\langle G\cdot v\rangle$. By the Bruit decomposition $G=BI_{1,K}\cup B\beta_{K}I_{1,K}$, we see
$\pi\subseteq\langle B\cdot v\rangle\subseteq\langle B\cdot w\rangle$.
Hence, we have proved $\pi=\langle B\cdot w\rangle$ for any $w\in\pi$, and the proposition follows. ∎
Remark 5.7.
The condition in Proposition 5.5 is merely sufficient: for the Steinberg representation $St$, its restriction to $B$ is irreducible, but $S_{+}$ does not annihilate the line $St^{I_{1,K}}$. More precisely, the representation $St$ is defined as $\textnormal{ind}^{G}_{B}1/(1)$, and its $I_{1,K}$-invariants is of one dimension. In the notation of section 3, the space $St^{I_{1,K}}$ is spanned by the image of $g_{1}$. By Proposition 3.5, we check that
$S_{+}g_{1}=-g_{2}$
which is just $S_{+}\overline{g_{1}}=\overline{g_{1}}$.
5.4 Proof of $(1)$ of Theorem 1.2
Theorem 5.8.
Let $\pi$ be a supersingular representation of $G$. Then $\pi\mid_{B}$ is irreducible.
Proof.
Let $w$ be any non-zero vector in $\pi$. We already know that (from the argument of Proposition 5.5)
$\pi^{I_{1,K}}\cap\langle B\cdot w\rangle\neq 0$.
Take any non-zero vector $v$ in the above space. As $\pi$ is supersingular, by Proposition 5.4 $v$ will be annihilated by $S^{k}_{+}$ for $k$ large enough. Let $m$ be the least positive integer satisfying that. Now the vector $v^{\prime}=S^{m-1}_{+}v$ is non-zero. By Proposition 3.2, $v^{\prime}$ is still $I_{1,K}$-invariant, and lies in $\langle B\cdot w\rangle$ by the form of $S_{+}$. It satisfies
$S_{+}v^{\prime}=0$.
We are done by Proposition 5.5. ∎
An immediate but interesting application of Theorem 5.8 is the following. Recall the Levi decomposition $B=H\ltimes N$, where $H$ is the diagonal subgroup and $N$ is the upper unipotent radical.
Corollary 5.9.
For a supersingular representation $\pi$ of $G$, we have
$\pi_{N}=0$,
i.e., the usual Jacquet module of $\pi$ (with respect to $B$) vanishes.
Proof.
Recall that $\pi_{N}=\pi/\pi(N)$, and the space $\pi(N):=\langle\pi(u)\cdot v-v\mid u\in N,v\in\pi\rangle$ is $B$-stable and non-zero. The assertion follows by Theorem 5.8. ∎
5.5 Proof of $(2)$ of Theorem 1.2
Theorem 5.10.
Give two smooth representations $\pi$ and $\pi^{\prime}$ of $G$. Suppose $\pi$ is supersingular. Then, we have
$\textnormal{Hom}_{G}(\pi,\pi^{\prime})\cong\textnormal{Hom}_{B}(\pi,\pi^{\prime}).$
Proof.
As we have one side inclusion $\textnormal{Hom}_{G}(\pi,\pi^{\prime})\hookrightarrow\textnormal{Hom}_{B}(\pi,\pi^{\prime})$, it suffices to prove any non-zero map $\phi$ in $\textnormal{Hom}_{B}(\pi,\pi^{\prime})$ is also a $G$-map.
We begin with a non-zero vector $v\in\pi^{I_{1,K}}$, and by replacing $v$ by a proper vector in the space $\langle I_{K}\cdot v\rangle$, we may assume $I_{K}$ acts on $v$ by a character $\chi$. As $\pi\mid_{B}$ is irreducible (Theorem 5.5), the map $\phi$ is injective whence $\phi(v)\neq 0$. As $\pi^{\prime}$ is smooth, $\phi(v)$ is fixed by $N^{\prime}_{m_{K}+2m}$ for some $m\geq 0$.
We assume $m\geq 1$, and proceed to find a non-zero vector $w\in\pi^{I_{1,K}}\cap\langle B\cdot v\rangle$ such that $\phi(w)$ is fixed by $N^{\prime}_{m_{K}+2m-2}$.
Case I: $S_{+}v\neq 0$. We take $w$ as $S_{+}v$, so it lies in $\pi^{I_{1,K}}\cap\langle B\cdot v\rangle$. Note that $I_{K}$ acts on $w$ still by the character $\chi$. Using Lemma 3.3, we see the vector $\phi(w)=S_{+}\phi(v)$ is fixed by $N^{\prime}_{m_{K}+2m-2}$.
Case II: $S_{+}v=0$ and $S_{-}v\neq 0$. We take $S_{-}v$ as $w$, which lies in $\pi^{I_{1,K}}$ (Lemma 3.2). Our assumption $S_{+}v=0$ means
$S_{K}w=0$,
which is equivalent to
$w=-\sum_{u\in N_{n_{K}}\setminus N_{n_{K}+1}/N_{n_{K}+2}}\beta_{K}u\alpha^{-1}v$
An application of (1) gives that for $u=n(\ast,\varpi^{n_{K}}_{E}t)\in N_{n_{K}}\setminus N_{n_{K}+1}$
$\beta_{K}u\alpha^{-1}=n(\ast,\varpi^{n_{K}}_{E}t^{-1})\alpha^{-1}h(\bar{t}^{-1})n^{\prime}(\ast,\varpi^{m_{K}+1}_{E}t^{-1})$,
hence we have $\beta_{K}u\alpha^{-1}\cdot v=\chi(h(\bar{t}^{-1}))n(\ast,\varpi^{n_{K}}_{E}t^{-1})\alpha^{-1}\cdot v$. In all we get
$w=-\sum_{u\in N_{n_{K}}\setminus N_{n_{K}+1}/N_{n_{K}+2}}\chi(h(\bar{t}^{-1}))u\alpha^{-1}\cdot v\in\langle B\cdot v\rangle$,
whence that
$\phi(w)=-\sum_{u\in N_{n_{K}}\setminus N_{n_{K}+1}/N_{n_{K}+2}}\chi(h(\bar{t}^{-1}))u\alpha^{-1}\cdot\phi(v)$
This, combined with Lemma 3.3 (especially (3)) and Remark 3.4, shows $\phi(w)$ is fixed by $N^{\prime}_{m_{K}+2m-2}$.
Case III: $S_{-}v=0$ (whence $S_{+}v=0$). By the definition of $S_{-}$, this assumption is same as $\sum_{u\in N_{n_{K}+1}/N_{n_{K}+2}}u\alpha^{-1}\cdot v=0$. Then by almost the same argument we have
$\beta_{K}v=-\sum_{u\in N_{n_{K}+1}\setminus N_{n_{K}+1}/N_{n_{K}+2}}\chi(h(\bar{t}^{-1}))u\alpha^{-1}\cdot v$.
Thus
$\phi(\beta_{K}v)=-\sum_{u\in N_{n_{K}+1}\setminus N_{n_{K}+1}/N_{n_{K}+2}}\chi(h(\bar{t}^{-1}))u\alpha^{-1}\cdot\phi(v)$
We conclude that $\phi(\beta_{K}v)$ is fixed by $N^{\prime}_{m_{K}+2m-2}$, using Lemma 3.3 (especially (3)) and Remark 3.4. Based on this, one verifies further that, for any $u\in N_{n_{K}}/N_{n_{K}+1}$, the vector $\phi(u\beta_{K}v)$ is still fixed by $N^{\prime}_{m_{K}+2m-2}$ , using Lemma 3.3 again.
$\bullet~{}S_{K}v=\sum_{u\in N_{n_{K}}/N_{n_{K}+1}}u\beta_{K}v\neq 0$. In this case, we take $S_{K}v$ as $w$. As
$\phi(w)=\sum_{u\in N_{n_{K}}/N_{n_{K}+1}}\phi(u\beta_{K}v)$,
we conclude that $\phi(w)$ is fixed by $N^{\prime}_{m_{K}+2m-2}$ by the proceeding remarks.
$\bullet~{}S_{K}v=\sum_{u\in N_{n_{K}}/N_{n_{K}+1}}u\beta_{K}v=0$. This assumption gives us that
$v\in\langle u\beta_{K}v|u\in N_{n_{K}}\setminus N_{n_{K}+1}/N_{n_{K}+1}\rangle$.
Explicitly, under this assumption we have
$v=-\sum_{u\in N_{n_{K}}\setminus N_{n_{K}+1}/N_{n_{K}+1}}\chi(h(\bar{t}^{-1}))u\beta_{K}\cdot v$
Consider the $K$-representation $\tau=\langle K\cdot v\rangle$. It is spanned by the set $\{v,u\beta_{K}v\mid u\in N_{n_{K}}/N_{n_{K}+1}\}$. Then we see it can be spanned by the set
$\{u\beta_{K}v\mid u\in N_{n_{K}}\setminus N_{n_{K}+1}/N_{n_{K}+1}\}$
Now take a weight $\sigma$ contained in $\tau$, and the unique line $\sigma^{I_{1,K}}$ is given by a non-zero vector which we take as our $w$. Then $w$ is a linear combination of the elements from the above set. This implies, as before that $\phi(w)$ is fixed by $N^{\prime}_{m_{K}+2m-2}$.
Note that in each case the group $I_{K}$ acts as a character on the vector $w$ we find. By repeating the process, we find a non-zero $w\in\pi^{I_{1,K}}$ on which $I_{K}$ acts as a character, such that $\phi(w)$ is fixed by $N^{\prime}_{m_{K}}$. Note that $\phi(w)$ is automatically fixed by $B\cap I_{1,K}$, whence the vector $\phi(w)$ lies in $(\pi^{\prime})^{I_{1,K}}$.
We proceed to complete the proof. Recall our condition that $\pi$ is supersingular. We apply Proposition 5.4 to the vector $w$. We find some $k\geq 1$ such that $S^{k}_{+}w=0$ but $S^{k-1}_{+}w\neq 0$. We write the vector $S^{k-1}_{+}w$ as $w^{\prime}$. Then $w^{\prime}$ is a non-zero $I_{1,K}$-invariant such that $S_{+}w^{\prime}=0$. This, as we already recorded in (7), gives that
$\beta_{K}w^{\prime}=-\sum_{u}u\alpha^{-2}h(\bar{t}^{-1})w^{\prime}-\sum_{u}u\alpha^{-1}h(\bar{t}^{-1})w^{\prime}$
where, in the first sum $u=n(\ast,\varpi^{n_{K}+2}_{E}t)$ goes through $(N_{n_{K}+2}\setminus N_{n_{K}+3})/N_{n_{K}+4}$; in the second sum $u=n(\ast,\varpi^{n_{K}+1}_{E}t)$ goes through $(N_{n_{K}+1}\setminus N_{n_{K}+2})/N_{n_{K}+2}$. As $\phi$ is a $B$-map, we firstly see that
$\phi(\beta_{K}w^{\prime})=-\sum_{u}u\alpha^{-2}h(\bar{t}^{-1})\phi(w^{\prime})-\sum_{u}u\alpha^{-1}h(\bar{t}^{-1})\phi(w^{\prime})$
But we also have $\phi(S_{+}w^{\prime})=S_{+}\phi(w^{\prime})=0$, which gives us similarly that
$\beta_{K}\phi(w^{\prime})=-\sum_{u}u\alpha^{-2}h(\bar{t}^{-1})\phi(w^{\prime})-\sum_{u}u\alpha^{-1}h(\bar{t}^{-1})\phi(w^{\prime})$
We conclude
$\phi(\beta_{K}w^{\prime})=\beta_{K}\phi(w^{\prime})$.
Recall that the vector $w^{\prime}$ and its image $\phi(w^{\prime})$ are both $I_{1,K}$-invariant. As $G=BI_{1,K}\cup B\beta_{K}I_{1,K}$ and the non-zero vector $w^{\prime}$ generates the representation $\pi$, we conclude that from the above equality that the $B$-map $\phi$ is indeed a $G$-map. ∎
Remark 5.11.
In early versions of this paper, we didn’t find a full proof of Theorem 5.10. We came to the above proof recently. The operator $S_{-}$ is an analogue of the element $\Pi=\begin{pmatrix}0&1\\ \varpi_{F}&0\end{pmatrix}$ in $GL_{2}(F)$, but the situation that $S_{-}v=0$ for some $v\in\pi^{I_{1,K}}$ can always happen, which is not the case for $GL_{2}(F)$. This is dealt with in our above Case III argument, and we consider it as a novelty of our work. It would be interesting to see whether it can be adapted to other problems.
Acknowledgements
[…]